OpenStudy (anonymous):

Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 80 miles per hour and train b is traveling at 96 miles per hour. Train A passes a station at 1:10 a.m. if train B passes the same station at 1:25 a.m., at what time will train b catch up to train A.

OpenStudy (anonymous):

1) difference in time between trains: 1:25 - 1:10 = 15 min (or 1/5 of an hour); 2) train B will travel 96/4 = 24 miles in 1/4 hour; 3) 24 = V * t = 16 * t 4) t=24/16 = 3/2 hour to catch train A 5) 1:25 + 1.5 -- 2:55

OpenStudy (anonymous):

There are not 75 minutes in an hour.